OpenStudy (haleyelizabeth2017):

The sum of two reciprocals of two consecutive positive integers is 17/72. Write an equation that can be used to find the two integers. What are the integers? *I found them, I just need it checked*

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

did you write something like \[\frac{1}{n}+\frac{1}{n+1}=\frac{17}{72}\]?

OpenStudy (haleyelizabeth2017):

\[\frac{1}{x}+\frac{1}{(x+1)}=\frac{17}{72}\] Yes lol

OpenStudy (misty1212):

kk you win

OpenStudy (haleyelizabeth2017):

I'll skip to the quad form....one sec

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

\[\frac{ 1 }{ x }+\frac{ 1 }{ x+1 }=\frac{ 17 }{ 72 },x \neq 0,x \neq-1\]

OpenStudy (misty1212):

just guess you will get it instantly

OpenStudy (misty1212):

\[8\times 9=72\\ 8+9=17\] that is all

OpenStudy (haleyelizabeth2017):

\[x=\frac{ -(-127) \pm \sqrt {(-127)^2-4(17)(-72)} }{ 2(17) }\] I can't guess...I have to show my work :/

OpenStudy (haleyelizabeth2017):

But I got a negative for the second number and I don't know what I did wrong :/

OpenStudy (haleyelizabeth2017):

After I did that, (I'm skipping some steps to save time): \[\frac {127 \pm 145}{34}\] x=8 x=-9/17 I'm not sure where I went wrong....

OpenStudy (anonymous):

a quadratic gives always two values. x=8 is integer ,hence is correct so consecutive integers are x,x+1 ?

OpenStudy (anonymous):

\[\frac{ x+1+x }{ x(x+1) }=\frac{ 17 }{ 72 }\] \[\frac{ 2x+1 }{ x^2+x }=\frac{ 17 }{ 72 }\] \(17 x^2+17x=144x+72\) \(17x^2-127x-72=0\) \[x=\frac{ 127\pm \sqrt{(-127)^2-4*17*72} }{ 2*17 }\]\[x=\frac{ 127\pm \sqrt{16129+4896} }{ 34 }=\frac{ 127\pm \sqrt{21025} }{ 34 }\] \[=\frac{ 127\pm145 }{ 34 }\] so you are correct .

OpenStudy (anonymous):

correction -4*17*-72=+4896

OpenStudy (haleyelizabeth2017):

Oh! Okay...thank you!